# FAST: Federated Active Learning with Foundation Models for Communication-efficient Sampling and Training

### v2

@@ IV-A Experimental Configuration
Datasets. We evaluate our method primarily on image classification tasks spanning both natural and medical benchmark datasets. Specifically, we use four natural image datasets—CIFAR10/100 [50], Tiny-ImageNet [51], and SVHN [52]—as well as two medical image datasets—PathMNIST and DermaMNIST [53]. To account for the inherent heterogeneity among clients, we consider three data distribution settings: IID, Non-IID, and heterogeneous inter-class diversity (i.e., variations in local class distributions) [7]. As shown in Table I, we report the total labeling budget and the corresponding number of training rounds. Following existing FAL approaches, each client trains its local model from scratch and iteratively selects 5% of the total dataset for annotation in each AL round, until reaching a predefined global labeling budget. To ensure fairness in labeling costs, we assume this global labeling budget is evenly shared among clients, such that each client queries the same number of samples per AL round.
@@
Baselines. We compare FAST with nine standard active learning (AL) strategies: Random, Entropy [32], Coreset [39], BADGE [54], LL4AL [37], GCNAL [41], and ALFA-Mix [55]. Although originally designed for centralized AL, these strategies can be independently applied on either a global or a local model within a federated environment. In our experiments, we employ the global model as the query selector for active sampling. We further include two federated AL (FAL) strategies, KAFAL [6] and LoGo [7]. For the Non-IID experiment in Figure 2, we select KAFAL as it is specifically tailored for global heterogeneity problems. Regarding the experiment on local heterogeneity in Table I, we add LoGo as the baseline considering its focus on solving heterogeneous data from the client level.
@@
Implementation Settings. We implement our proposed FAST method in PyTorch using the Flower FL framework [56]. As our primary federated learning (FL) strategy, we adopt FedAvg [1], and additionally evaluate on FedProx [2] and FedNova [57] to examine the robustness of FAST across different FL paradigms (see Table III). Our experiments primarily target cross-silo settings with full client participation, involving a total of $10$ clients. Each AL round spans $T=100$ federated communication rounds, and each client executes $\tau=5$ local stochastic gradient descent (SGD) steps per round. In alignment with prior work [7, 6], we employ a four-layer CNN as our main model architecture and employ a ResNet-8 network for ablation studies on communication efficiency. We simulate the Non-IID data partitions by sampling from a Dirichlet distribution with a concentration parameter of $\alpha=0.1$, where smaller values of $\alpha$ indicate greater data heterogeneity across clients [58]. For the implementation of FAST, we initialize with 1% of labeled data and employ a frozen SigLIP [11] as the foundation model for feature extraction and weak labeling in the two-pass process.
@@ IV-B Results
Comparison on heterogeneous inter-class diversity data. We first evaluate the performance of FAST in comparison with other baseline methods on datasets characterized by high levels of local heterogeneity. In this context, each client shares the same pool of classes but exhibits varying inter-class distributions. Conventional AL methods aim to minimize labeling efforts by selecting a small subset of instances based on their informativeness across the entire dataset. However, in a decentralized setting where each local dataset maintains distinct class distributions, such imbalanced data partitions often lead to inconsistent knowledge sharing.
Consequently, the selected samples may not be representative or sufficiently informative for all clients, thereby hindering the overall learning performance. As presented in Table I, we compare the performance of FAST with other existing AL strategies under a one-shot setting, wherein only a single active sampling round is conducted. We observe that FAST outperforms all baseline methods even within the constraints of this one-shot scenario. Notably, in this experiment, the server exhausts 5% of the labeling budget per round until reaching the total budget limit.
@@ IV-C Ablation Studies
Ablation on different federated learning strategies. We investigate the impact of various FL strategies on the performance of FAST under a fixed labeling budget of 20%. Table III reports the accuracy across five benchmark datasets. Notably, FedNova offers marginal yet consistent improvements over FedAvg and FedProx on most datasets, indicating that FAST is compatible with advanced FL aggregation strategies and can further support heterogeneous scenarios. These findings confirm the robustness of FAST under different federated aggregation schemes.
TABLE III: Performance of Our Method Across Different Federated Learning Strategies with 20% Labeling Budget
@@
Effect of Different Foundation Model Selections on FAST. We next evaluate how the choice of foundation model for the preliminary pass in FAST influences its overall performance. Specifically, we compare three vision-language models—CLIP, EvaCLIP, and SigLIP—along with an image-specific model, DINOv2, using a pre-trained ResNet-50 as the baseline. As shown in Table IV, EvaCLIP consistently achieves the highest accuracy across all datasets, followed closely by SigLIP and DINOv2. This underscores the importance of rich representation knowledge for enhancing weak labeling quality in the preliminary pass. Furthermore, the results suggest that leveraging expressive embeddings can significantly improve active sampling outcomes, even under constrained annotation budgets.
TABLE IV: Performance Comparison of Our Methods with Different Foundation Models
@@
TABLE V: Performance of Our Method with Varying Labeling Budgets. Training with the FedAvg strategy using a CNN-4 model, 10 clients, 100 rounds.
@@
Ablation on Labeling Budget. To assess the scalability of FAST with respect to the labeling budget, we evaluate its performance under varying labeling budgets ranging from 0% to 80%. Table V illustrates the accuracy of FAST across various datasets as the labeling budget increases. The results demonstrate a positive correlation between the labeling budget and model accuracy, with significant performance improvements observed as the budget increases. For instance, on CIFAR-10, accuracy improves from 75.92% at 0% budget to 77.48% at 80% budget. Similar trends are observed across CIFAR-100 and Path-MNIST, indicating the effectiveness of FAST in leveraging additional unlabeled data to enhance model performance under constrained labeling budgets.
@@ V Conclusion
In this paper, we introduced a two-pass FAL framework, FAST, designed to address the critical challenges of limited annotation budgets and communication-intensive sampling processes in FAL. Our approach leverages robust representation-based knowledge from foundation models to efficiently query informative unlabeled data for annotation, thereby minimizing human effort and communication overhead. Extensive experiments on diverse vision datasets demonstrate that FAST consistently outperforms existing FAL methods in terms of both predictive performance and communication cost. These findings underscore the potential of leveraging foundation models to enhance FAL under realistic resource constraints. Future directions include exploring more sophisticated query strategies within FAST and quantifying weak labeling quality, enabling label correction prior to final human annotation and thereby further enhancing performance and communication efficiency.
@@ Acknowledgments
TABLE VI: Performance of uncertainty sampling strategies on weak-labeled data across various datasets. Training with 10 clients for 100 rounds, utilizing a 4-layer CNN network. Evaluating with FedAvg.
@@
We evaluate several uncertainty-based query strategies: norm-based, entropy-based, least confidence, smallest margin, and largest margin—applied to the prototype vectors computed for each weakly labeled sample. As summarized in Table VI, the results are generally comparable across different datasets, suggesting that the prototype-based logits capture the key uncertainty information leveraged by a variety of query strategies. This underscores the effectiveness of the prototype representation in identifying highly uncertain samples for human refinement.

### v3

@@ IV-A Experimental Configuration
Datasets. We evaluate our method primarily on image classification tasks spanning both natural and medical benchmark datasets. Specifically, we use four natural image datasets—CIFAR10/100 [50], Tiny-ImageNet [51], and SVHN [52]—as well as two medical image datasets—PathMNIST and DermaMNIST [53]. To account for the inherent heterogeneity among clients, we consider three data distribution settings: IID, Non-IID, and heterogeneous inter-class diversity (i.e., variations in local class distributions) [7]. As shown in Table I, we report the total labeling budget and the corresponding number of training rounds. Following existing FAL approaches, each client trains its local model from scratch and iteratively selects 5% of the total dataset for annotation in each AL round, until reaching a predefined global labeling budget. We use a 20% labeling budget for most of our experiments and ablation studies. To ensure fairness in labeling costs, we assume this global labeling budget is evenly shared among clients, such that each client queries the same number of samples per AL round.
@@
Baselines. We compare FAST with nine standard active learning (AL) strategies: Random, Entropy [32], Coreset [39], BADGE [54], LL4AL [37], GCNAL [41], and ALFA-Mix [55]. Although originally designed for centralized AL, these strategies can be independently applied on either a global or a local model within a federated environment. In our experiments, we employ the global model as the query selector for active sampling. We further include two federated AL (FAL) strategies, KAFAL [6] and LoGo [7]. For the Non-IID experiment in Figure 2, we select KAFAL as it is specifically tailored for global heterogeneity problems. Regarding the experiment on IID datasets in Table I, we add LoGo as the baseline considering its focus on solving heterogeneous data from the client level.
@@
Implementation Settings. We implement our proposed FAST method in PyTorch using the Flower FL framework [56]. As our primary federated learning (FL) strategy, we adopt FedAvg [1], and additionally evaluate on FedProx [2] and FedNova [57] to examine the robustness of FAST across different FL paradigms (see Table IV). Our experiments primarily target cross-silo settings with full client participation, involving a total of $10$ clients. Each AL round spans $T=100$ federated communication rounds, and each client executes $\tau=5$ local stochastic gradient descent (SGD) steps per round. In alignment with prior work [7, 6], we employ a four-layer CNN as our main model architecture and employ a ResNet-8 network for ablation studies on communication efficiency. We simulate the Non-IID data partitions by sampling from a Dirichlet distribution with a concentration parameter of $\alpha=0.1$, where smaller values of $\alpha$ indicate greater data heterogeneity across clients [58]. For the implementation of FAST, we initialize with 1% of labeled data and employ a frozen SigLIP [11] as the foundation model for feature extraction and weak labeling in the two-pass process. We conducted all experiments on 2 NVIDIA A10 GPUs with Intel Xeon Gold 6342 CPUs (2.80GHz).
@@ IV-B Results
Comparison on heterogeneous inter-class diversity data. We first evaluate the performance of FAST in comparison with other baseline methods on datasets characterized by high levels of local heterogeneity. In this context, each client shares the same pool of classes but exhibits varying inter-class distributions. We deliberately fix FAST’s labeling budget at 5% to underscore its cost-effectiveness (Table I). With just one AL round and 5% labeled data, FAST matches or exceeds the accuracy of baselines that consume up to 40% budget over multiple AL rounds. In contrast, conventional AL methods sample 5% per round iteratively until they exhaust the same budget, incurring additional communication and computation. FAST achieves competitive performance through a single, one-shot active labeling round. This design choice highlights FAST’s significant efficiency advantages in reducing communication overhead.
Conventional AL methods aim to minimize labeling efforts by selecting a small subset of instances based on their informativeness across the entire dataset. However, in a decentralized setting where each local dataset maintains distinct class distributions, such imbalanced data partitions often lead to inconsistent knowledge sharing. Consequently, the selected samples may not be representative or sufficiently informative for all clients, thereby hindering the overall learning performance. As presented in Table I, we compare the performance of FAST with other existing AL strategies under a one-shot setting, wherein only a single active sampling round is conducted. We observe that FAST outperforms all baseline methods even within the constraints of this one-shot scenario. Notably, in this experiment, the server exhausts 5% of the labeling budget per round until reaching the total budget limit.
@@
Moreover, our analysis in Table II decomposes the FAST framework into distinct components for qualitative evaluation. Notably, we observe that applying uncertainty sampling ( 1+ 2+ 3) consistently outperforms random sampling in the refinement pass ( 1+ 2+ 4) across various datasets. These results empirically validate the effectiveness of uncertainty sampling within the second pass of our framework.
@@ IV-C Ablation Studies
Analysis on computational and communications overhead in FAST. In FL, communication overhead is most commonly measured either by the number of rounds or by the total volume of parameter transmissions (i.e., model uploads and downloads) [59, 60, 61]. As shown in Table I and Figure 2 of the main paper, FAST achieves higher accuracy within 100 FL rounds while reducing the number of communication rounds by a factor of eight. Table III reports the aggregate parameter‐transmission cost of each method. In FAST’s preliminary stage, a small set of labeled embeddings is uploaded to and redistributed from the server for weak labeling, incurring an initial communication cost. However, by leveraging representation‐based labeling and minimal human annotation, FAST greatly reduces the required number of FL rounds and thus the overall communication volume. In contrast, conventional FAL methods incur identical model‐aggregation and distribution costs in every round, and their extensive active‐learning cycles result in comparable or higher total communication overhead.
Furthermore, we compare computational costs in terms of wall‐clock time (Table III). The only additional expense in FAST’s preliminary stage is encoding local data via the foundation model’s encoder for uncertainty estimation. By contrast, LoGo [7] performs both macro‐ and micro‐level informativeness evaluations using local and global models in each active‐learning round, imposing a substantially higher runtime burden. As the table shows, on CIFAR-10 LoGo incurs approximately 2310 s more wall‐clock time than random sampling over the same number of rounds.
TABLE III: Comparison of classification accuracy (%), communication cost (MB), and wall-clock time (s) for FAST, LoGo, and random sampling on CIFAR-10, SVHN, and PathMNIST under the IID setting. We apply FedAvg on a four-layer CNN classifier (with a size of 0.45 MB on the CIFAR-10 dataset). FAST is evaluated with one AL round (100 FL rounds), whereas Random and LoGo use 8 AL rounds (i.e., 800 FL rounds).
Method CIFAR-10 CIFAR-100 SVHN
Acc. (%) Comm. Cost (MB) Walltime (s) Acc. (%) Comm. Cost (MB) Walltime (s) Acc. (%) Comm. Cost (MB) Walltime (s)
Random 69.14 7090.94 30398.95 32.67 8502.69 37258.75 85.47 6969.79 58064.42
LoGo 71.92 7090.94 32709.14 34.27 8502.69 39268.24 87.08 6969.79 61638.97
Ours    77.16 $\uparrow$        902.54 (87.27%) $\downarrow$      7342.52 (76.73%) $\downarrow$    41.94 $\uparrow$    1079.56 (87.30%) $\downarrow$     15104.38 (60.53%) $\downarrow$    88.79 $\uparrow$     896.72 (87.13%) $\downarrow$    19178.83 (67.96%) $\downarrow$
Ablation on different federated learning strategies. We investigate the impact of various FL strategies on the performance of FAST under a fixed labeling budget of 20%. Table IV reports the accuracy across five benchmark datasets. Notably, FedNova offers marginal yet consistent improvements over FedAvg and FedProx on most datasets, indicating that FAST is compatible with advanced FL aggregation strategies and can further support heterogeneous scenarios. These findings confirm the robustness of FAST under different federated aggregation schemes.
TABLE IV: Performance of Our Method Across Different Federated Learning Strategies with 20% Labeling Budget
@@
Effect of Different Foundation Model Selections on FAST. We next evaluate how the choice of foundation model for the preliminary pass in FAST influences its overall performance. Specifically, we compare three vision-language models—CLIP, EvaCLIP, and SigLIP—along with an image-specific model, DINOv2, using a pre-trained ResNet-50 as the baseline. As shown in Table V, EvaCLIP consistently achieves the highest accuracy across all datasets, followed closely by SigLIP and DINOv2. This underscores the importance of rich representation knowledge for enhancing weak labeling quality in the preliminary pass. Furthermore, the results suggest that leveraging expressive embeddings can significantly improve active sampling outcomes, even under constrained annotation budgets.
TABLE V: Performance Comparison of Our Methods with Different Foundation Models
@@
TABLE VI: Performance of Our Method with Varying Labeling Budgets. Training with the FedAvg strategy using a CNN-4 model, 10 clients, 100 rounds.
@@
Ablation on Labeling Budget. To assess the scalability of FAST with respect to the labeling budget, we evaluate its performance under varying labeling budgets ranging from 0% to 80%. Table VI illustrates the accuracy of FAST across various datasets as the labeling budget increases. The results demonstrate a positive correlation between the labeling budget and model accuracy, with significant performance improvements observed as the budget increases. For instance, on CIFAR-10, accuracy improves from 75.92% at 0% budget to 77.48% at 80% budget. Similar trends are observed across CIFAR-100 and Path-MNIST, indicating the effectiveness of FAST in leveraging additional unlabeled data to enhance model performance under constrained labeling budgets.
@@ V Conclusion
In this paper, we introduced a two-pass FAL framework, FAST, designed to address the critical challenges of limited annotation budgets and communication-intensive sampling processes in FAL. Our approach leverages robust representation-based knowledge from foundation models to efficiently query informative unlabeled data for annotation, thereby minimizing human effort and communication overhead. Extensive experiments on diverse vision datasets demonstrate that FAST consistently outperforms existing FAL methods in terms of both predictive performance and communication cost. These findings underscore the potential of leveraging foundation models to enhance FAL under realistic resource constraints. Future directions include exploring more sophisticated query strategies within FAST and quantifying weak labeling quality. Developing an additional filtering mechanism after the two-pass labeling process to enable label correction prior to final human annotation, thereby further enhancing performance and communication efficiency.
@@ Acknowledgments
TABLE VII: Performance of uncertainty sampling strategies on weak-labeled data across various datasets. Training with 10 clients for 100 rounds, utilizing a 4-layer CNN network. Evaluating with FedAvg.
@@
We evaluate several uncertainty-based query strategies: norm-based, entropy-based, least confidence, smallest margin, and largest margin—applied to the prototype vectors computed for each weakly labeled sample. As summarized in Table VII, the results are generally comparable across different datasets, suggesting that the prototype-based logits capture the key uncertainty information leveraged by a variety of query strategies. This underscores the effectiveness of the prototype representation in identifying highly uncertain samples for human refinement.